OpenStudy (anonymous):

Please help with my math questions!!! What is the value of expression? 5^-8 * 5^-4 A. 5^32 B. 5^4 C. 5^−4 D. 5^−12 What is the missing value? 7 (empty box) ----------- 7^5 = 7^-2 A. 7 B. 3 C. −3 D. −7 What is the missing exponent? (2^-5) [please tell number here] = 2^-15 what is the value of the expression? 9^8 -------- 9^-4 * 9^10 A. 9^−48 B. 9^−2 C. 9^2 D. 9^22 what is the missing exponent? 5^8 ------ {5^-4)^-3 = 5[please type number here]

OpenStudy (anonymous):

I got the first question on my own but I need help with the rest

OpenStudy (anonymous):

@zimmah

OpenStudy (anonymous):

\[\Large \frac{ 7^? }{ 7^5 } = 7^{-2}\]

OpenStudy (anonymous):

or in other words \(\Large 7^{?-5} = 7^{-2}\)

OpenStudy (anonymous):

or better yet, x-5 = -2

OpenStudy (anonymous):

no no no no, I got that already. I need help with 3, 4 and 5.

OpenStudy (anonymous):

oh right, i thought you only had the first one

OpenStudy (anonymous):

Sorry, forgot to add it :)

OpenStudy (anonymous):

\[\Large (2^{-5})^? = 2^{-15}\]

OpenStudy (anonymous):

that's how i read the question

OpenStudy (anonymous):

I put -10 is that right? where the question mark is

OpenStudy (anonymous):

no that would make a much larger number

OpenStudy (anonymous):

you are not multiplying them, you are putting them the power of each other. The rule for that becomes \(\Large (a^n)^q =a^{n*q}\)

OpenStudy (anonymous):

it's like (2^3)^2 is like (2*2*2)^2 so that would be (2*2*2)(2*2*2) get it?

OpenStudy (anonymous):

uh I don't understand that though ;-; It says what is the missing expoent needed to make the = sign true.

OpenStudy (anonymous):

So putting -5 and -10 would make it -15 right?

OpenStudy (anonymous):

let's do it the other way around if we want to know 3*?=21 we solve it by 21/3=7 so if we want to know (2^-5) [please tell number here] = 2^-15 we do -15/-5 = to find which exponent we need

OpenStudy (anonymous):

no, that would make 50

OpenStudy (anonymous):

jeez then XD how can I make it so it's true, by the way, this is 8th work :P

OpenStudy (anonymous):

you would get this: \(\Large \frac{ 1 }{ 2^5*2^5*2^5*2^5*2^5*2^5*2^5*2^5*2^5*2^5 }\)

OpenStudy (anonymous):

oh actually you have to divide that by one

OpenStudy (anonymous):

so you'd get 2^50

OpenStudy (anonymous):

you are confusing \(\Large (a^n)^q =a^{n*q} \) with \(\Large a^n*a^q=a^{n+q} \)

OpenStudy (anonymous):

note the difference

OpenStudy (anonymous):

it's easy to get these mixed up, but they are fundamentally different

OpenStudy (anonymous):

so if \(\Large 2^{?*-5} = 2^{-15} \) than what needs to be on the ? for the equation to be right?

OpenStudy (anonymous):

@mario0012 do you need any more explanation? i can work it out in detail if you want

OpenStudy (anonymous):

Thank you :)